OpenStudy (anonymous):

sin^2x-cos^2x/1-cot^2x

OpenStudy (anonymous):

Okay, I'm guessing it's actually (sin^2x-cos^2x)/(1-cot^2x)... First of all, let's write \[1-\cot^2x\] in terms of sin and cos: \[1 - \cot^2x =\] \[1 - (\cos^2x/\sin^2x) =\] \[(\sin^2x / \sin^2x) - (\cos^2x) / (\sin^2x) = \] \[(\sin^2x - \cos^2x) / \sin^2x\] Well, that's nifty, because now I've got something similar on the top and the bottom. \[(\sin^2x - \cos^2x) / ((\sin^2x - \cos^2x) / \sin^2x) = \] \[((\sin^2x - \cos^2x) / 1) * (\sin^2x / (\sin^2x - \cos^2x)) = \] \[1 / \sin^2x =\] \[\csc^2x\]